# Deep learning joint extremes of metocean variables using the SPAR model

### v1

@@ 3.2.3 Selecting an architecture
To select our ‘optimal’ architecture for the application detailed in Section 4, we perform a grid-search over architecture choices. For each configuration, we estimate a range of model fit diagnostics; these are discussed below. The optimal architecture is then chosen as the configuration which visually provides the best model diagnostics. We found that for both MLPs, a simple architecture is preferable: $L=3$ hidden layers, with $h_{l}=16$ neurons per layer. This results in two MLPs, each comprising approximately 650 estimable parameters; inference for these models is not computationally demanding, and can be conducted on a standard laptop.
As with univariate POT models, selecting a suitable threshold for our model is critical. Too low a threshold will result in the asymptotic arguments motivating the use of the GP model not being applicable, causing bias; whereas too high a threshold will result in few observations to fit to, resulting in higher variance. The process of threshold selection for our model is directly analogous to that in univariate problems. That is, we fit the model for a range of threshold exceedance probabilities $\zeta$, and check for stability of inferences and goodness of fit. The threshold is then selected as the largest value of $\zeta$ for which inferences are approximately stable for $\zeta_{0}<\zeta$. In our application, we found that $\zeta=0.1$ was suitable.
We remark that while our selected architecture works well for our application, we do not advocate the general use of these hyperparameters. Instead, we recommend that practitioners who apply our framework perform a similar grid-search, and and use post-fit diagnostics to select the optimal architecture.
@@ 4.1 Dataset
In this section, we consider the application of the SPAR model to a hindcast dataset consisting of 31-years of wind and wave variables from 01/01/1990 to 31/12/2020, for a site in the Celtic Sea, off the south-west coast of the UK, which has been identified for development of floating wind farm projects. The dataset consists of hourly values of significant wave height $(H_{s})$, mean wave period $(T_{m})$, mean wave direction $(\theta_{wave})$, hourly mean wind speed at 10 m above sea level $(U_{10})$, and wind direction $(\theta_{wind})$. These variables all influence the motion and loading of floating wind turbines, and understanding their joint extremes is important for design. Since directional variables are periodic, it does not make sense to talk about ‘extreme directions’. Instead, we work with the x- and y-components of wave height and wind speed, defined as $H_{x}=H_{s}\cos(\theta_{wave})$, $H_{y}=H_{s}\sin(\theta_{wave})$, $U_{x}=U_{10}\cos(\theta_{wind})$, and $U_{y}=U_{10}\sin(\theta_{wind})$.
Figure 2: Upper right plots: Scatter of pairwise relations between normalised variables for simulated (black) and observed (grey) data. Dashed red lines in plots of $(x_{3},x_{5})$ and $(x_{4},x_{5})$ are lines of constant wave steepness $s=0.08$. Plots on diagonal are QQ plots of simulated vs observed data. Lower left plots: Empirical density of pairs of angular components of observed data (yellow = high density, blue = low density).
@@ 4.2 Normalisation and choice of origin
As different physical variables have different scales, we normalise each variable by its standard deviation. If we did not do this, then angles would tend to be clustered in the plane of whichever variable has the largest scale (wind speed in this case). We also need to define an origin in order to transform to polar coordinates. In previous work using SPAR, [mackay2025spar, murphy2024inference] defined the origin at the mean of each variable. In our application, more care must be taken when defining an appropriate origin. For the model to provide a useful description of the extremes at all angles, the support of the density, $\mathrm{supp}(f_{\mathbf{X}}):=\{\mathbf{x}\in\mathbb{R}^{d}:f_{\mathbf{X}}(\mathbf{x})>0\}$, must be star-shaped with respect to the chosen origin [hansen2020]. That is, given an origin $\mathbf{x}_{0}\in\mathbb{R}^{d}$ and any point $\mathbf{x}\in\mathrm{supp}(f_{\mathbf{X}})$, the line segment from $\mathbf{x}_{0}$ to $\mathbf{x}$ is contained in $\mathrm{supp}(f_{\mathbf{X}})$. Under this assumption, all rays from the origin reach the ‘edges’ of the distribution without passing through regions which have zero density. In this way, the data-cloud has a well-defined ‘inside’ and ’outside’, with the ‘outer’ region considered ‘extreme’, and the representation of the radial component in this region by a GP distribution is reasonable. This assumption can be verified by checking plots of the density of observations along various rays from the origin, e.g., histograms of the observed radial variable within small angular ranges. However, as discussed further below, in higher dimensional spaces, a large number of angles are required to obtain a reasonable coverage of the surface of $\mathbb{S}^{d-1}$, and visual inspection of plots of the radial density over each angular range is time consuming. Ultimately a data-driven approach for selecting an optimal choice of origin would be best.
@@
In this example, we use some physical insights to define an appropriate origin. Firstly, wave breaking limits the maximum possible wave height for a given wave period, with the limit related to wave steepness given by $s=2\pi H_{s}/(gT_{m}^{2})$. In the three-dimensional subspace containing the variables $(H_{x},H_{y},T_{m})$, this results in a conical-shaped bound on the data, centred along the axis $H_{x}=H_{y}=0$, with the radius of the cone given by $H_{s,max}=s_{max}gT_{m}^{2}/(2\pi)$, where $s_{max}$ is the limiting steepness. The limit value of $s_{max}$ depends on the water depth and wind speed (among other factors), but the maximum value was found to be around $s_{max}\approx 0.08$ for our dataset. This conical shape to the distribution (see Figures 2 and 8) suggests that an appropriate choice of origin should be somewhere on the axis $H_{x}=H_{y}=0$. Due to the physical dependence of wave height on wind speed, we also locate the origin at $U_{x}=U_{y}=0$. The choice of origin for $L_{T}$ is somewhat arbitrary, but experimentation showed that using the mean value of $L_{T}$ gave satisfactory results.
@@
The angular and radial variables are therefore defined with respect to the normalised variables given by $X_{1}=U_{x}/\mathrm{STD}(U_{x})$, $X_{2}=U_{y}/\mathrm{STD}(U_{y})$, $X_{3}=H_{x}/\mathrm{STD}(H_{x})$, $X_{4}=H_{y}/\mathrm{STD}(H_{y})$, and $X_{5}=(L_{T}-\mathrm{mean}(L_{T}))/\mathrm{STD}(L_{T})$, where $\mathrm{STD}(\cdot)$ denotes the standard deviation function. The pairwise relations between these normalised variables are shown in Figure 2. A radial grid has been overlaid to illustrate that these two-dimensional projections are approximately star-shaped with respect to this origin. Lines of bounding steepness $s=0.08$ are shown in the plots of $(X_{3},X_{5})$ and $(X_{4},X_{5})$ as dashed lines. It can be seen that there is far less scatter in the variables close to these bounds due to the physical limitations. Another feature of the data that is evident is the strong positive correlation between the x- and y-components of the wave height and wind speed.
@@ 4.3 Exploratory data analysis
Before assessing the model fit, it is useful to consider various visualisations of the data, in order to understand how it is distributed over the five-dimensional space. The plots below the diagonal in Figure 2 show the empirical densities of pairs of angular components $(W_{i},W_{j})$. By definition, these variables must fall within the unit circle. Any large gaps in the observed values indicate that it is not possible to fit the model at these angles. The objective of using the MLP model of the angular variation of the radial distribution is to estimate a relatively smooth variation with angle. The model should therefore be able to smooth over small angular ranges with no observations. However, the model is unlikely to be able to accurately estimate the behaviour of the radial component over large angular ranges with little or no data.
@@
Figure 3: Empirical joint density of wind direction and wave direction.
To assess the variation in the density of angles around a circle, we could plot the number of observations within discrete angular ranges as a histogram. In higher dimensions, visualisation becomes more difficult, but a similar approach can be taken. We count the number of observations within a small angular range of a pseudo-regularly spaced grid of points on the sphere. It is not possible to define evenly-spaced points on the surface of a (hyper) sphere in three or more dimensions. To address this issue, we take the approach proposed in [mackay2023-DIFORM], and define a regular grid of points on the $L^{1}$ sphere and project this onto the surface of the $L^{2}$ sphere. This is computed by creating a regular grid of points in the cube $[-1,1]^{d}$, with $12m+$ 1 points along each dimension: $U_{grid}=\{(i_{1}/m,...,i_{d}/m):i_{j}\in\{-m,...,m\},j=1,...,d\}$. Then, we define $U_{1}=\{\mathbf{u}\in U_{grid}:\|\mathbf{u}\|_{1}=1\}$, where $\|\cdot\|_{1}$ is the $L^{1}$ norm given by $\|(u_{1},...,u_{d})\|_{1}=|u_{1}|+\cdots+|u_{d}|$. Finally, we compute a set of direction vectors $U\subset\mathbb{S}^{d-1}$ by $U=\{\mathbf{u}/\|\mathbf{u}\|_{2}:\mathbf{u}\in U_{1}\}$. This is illustrated in Figure 4 for the case $m=5$ and $d=3$.
The dot product of two unit vectors is the cosine of the angle between them. Therefore, for each $\mathbf{u}_{j}\in U$ we count the number of observed angles with $\arccos(\mathbf{w}_{i}\cdot\mathbf{u}_{j})<\theta_{max}$, where $\theta_{max}$ is some prescribed range. Although there may be some overlap between the ranges defined above, this analysis still gives an indication of how the density of angles varies over the sphere.
Figure 6 shows the empirical CDF of the number of observations in a cell chosen at random, for a cell radius of $\theta_{max}=15^{\circ}$ and a set of 1002 direction vectors (generated using $m=5$ in the definition of $U_{grid}$ above). One feature that is evident is that over 50% of angular cells contain no observations. This is due to the particular choice of origin, which was selected to meet the assumption of a star-shaped distribution. Figure 6 also shows the number of observations in each of the 32 orthants in $\mathbb{R}^{5}$ (an orthant is the $d$-dimensional analogue of a quadrant of the plane). There are five orthants which contain no observations, and a further five which contain fewer than 100 observations (out of a total of 271,704 observations). The SPAR model estimates of the extremes in these regions will therefore be highly uncertain in the orthants with little data. However, Figure 6 shows scatter plots of the maximum observed values of $H_{s}$ and $U_{10}$ in each of the 1002 angular cells against the corresponding number of observations in the cell, indicating that larger values of these variables tend to coincide with higher angular densities. The higher uncertainties associated with the lower occurrence regions should therefore have less effect on the global extremes.
@@
Figure 8 shows a scatter plot of $T_{m}$ against $H_{x}$ and $H_{y}$. (A plot of the normalised variables $(X_{3},X_{4},X_{5})$ would look similar, but the non-normalised variables are shown here to aid physical interpretation). The conical shaped bound imposed by the limiting wave steepness is evident. Another feature that is apparent is that the data cloud is hollow on the side $H_{x}<0$. This is because of fetch limitations in this direction, meaning that waves propagating towards the west are steeper wind-driven waves, so that $H_{s}$ and $T_{m}$ are strongly correlated in this region. This violates the assumption of the distribution being star-shaped with respect to the choice of origin. However, the ‘edge region’ that is not modelled corresponds to low values of $H_{s}$ at a given $T_{m}$ and wave direction, which is less critical for extreme responses.
For the present choice of origin at $(H_{x},H_{y},T_{m})=(0,0,6.2)$, it appears that there are some sharp changes in the distribution with the vertical angle, for rays towards the negative x direction. This might be hard to capture with an ANN. Nevertheless, we can assess how well the model performs, given these challenges.
@@
Figure 7: Scatter plot of observations of x-y components of $H_{s}$ and $T_{m}$. Colour indicates value of $H_{s}=\sqrt{H_{x}^{2}+H_{y}^{2}}$ (blue=low, yellow=high).
Figure 8: Scatter plots of pairwise relations between simulated (black) and observed (grey) data in original variable space.
4.4 Diagnostic plots
Creating diagnostic plots to assess goodness-of-fit is challenging in multivariate problems. Assessing pairwise relations, as described above, is one option, where observed and simulated data can be overlaid to visually assess the plausibility of the fitted model, as shown in Figure 2, where the simulated sample from the fitted SPAR model is of equal size to the observed sample. Overall, the simulated relationships between variables closely follow those of the observations. Figure 8 shows scatter plots of the data transformed back to the original scale, to assess the modelled relationship between $H_{s}$, $T_{m},$ and $U_{10}$. Again, the simulated data is a good representation of the observations. The model predicts some larger values of steep waves than observed. However, these occur at lower wave heights, so are less critical for design purposes.

### v2

@@ 3.2.3 Selecting an architecture
To select our ‘optimal’ architecture for the application detailed in Section 4, we perform a grid-search over architecture choices. For each configuration, we estimate a range of model fit diagnostics; these are discussed below. The optimal architecture is then chosen as the configuration which visually provides the best model diagnostics. We found that, for both MLP models, a simple architecture is preferable: $L=3$ hidden layers, with $h_{l}=16$ neurons per layer. This results in two MLPs, each comprising approximately 650 estimable parameters; inference for these models is not computationally demanding, and can be conducted on a standard laptop.
As with univariate POT models, selecting a suitable threshold for our model is critical. Too low a threshold will result in the asymptotic arguments motivating the use of the GP model not being applicable, causing bias; whereas too high a threshold will result in too few observations for fitting, resulting in higher variance. The process of threshold selection for our model is directly analogous to that in univariate problems. That is, we fit the model for a range of threshold exceedance probabilities $\zeta$, and check for stability of inferences and goodness of fit. The threshold is then selected as the largest value of $\zeta$ for which inferences are approximately stable for $\zeta_{0}<\zeta$. In our application, we found that $\zeta=0.1$ was suitable.
We remark that while our selected architecture works well for our application, we do not advocate the general use of these hyperparameters. Instead, we recommend that practitioners who apply our framework perform a similar grid-search, and use post-fit diagnostics to select the optimal architecture.
@@ 4.1 Dataset
In this section, we consider the application of the SPAR model to a hindcast dataset consisting of 31-years of wind and wave variables from 01/01/1990 to 31/12/2020, for a site in the Celtic Sea, off the south-west coast of the UK, which has been identified for development of floating wind farm projects. The dataset consists of hourly values of significant wave height $(H_{s})$, mean wave period $(T_{m})$, mean wave direction $(\theta_{wave})$, hourly mean wind speed at 10 m above sea level $(U_{10})$, and wind direction $(\theta_{wind})$. Both wind and wave directions are defined as the directions in which the winds or waves are ‘going to’. These variables all influence the motion and loading of floating wind turbines, and understanding their joint extremes is important for design. Since directional variables are periodic, it does not make sense to talk about ‘extreme directions’. Instead, we work with the x- and y-components of wave height and wind speed, defined as $H_{x}=H_{s}\cos(\theta_{wave})$, $H_{y}=H_{s}\sin(\theta_{wave})$, $U_{x}=U_{10}\cos(\theta_{wind})$, and $U_{y}=U_{10}\sin(\theta_{wind})$.
Figure 3: Upper right plots: Empirical joint densities of pairs of normalised observations. Dashed red lines in plots of $(x_{3},x_{5})$ and $(x_{4},x_{5})$ are lines of constant wave steepness $s=0.08$. Plots on diagonal are histograms of each normalised variable. Lower left plots: Empirical density of pairs of angular components of observed data (yellow = high density, blue = low density). The normalised variables $(X_{1},...,X_{5})$ correspond to normalised $(U_{x},U_{y},H_{x},H_{y},\log(T_{m}))$.
@@ 4.2 Normalisation and choice of origin
As different physical variables have different scales, we normalise each variable by its standard deviation. Without this, angles would tend to be clustered in the plane of whichever variable has the largest scale (wind speed in this case). We also need to define an origin in order to transform to polar coordinates. In previous work using SPAR [mackay2025spar, murphy2024inference], the origin was defined at the mean of each variable. In our application, more care must be taken when defining an appropriate origin. For the model to provide a useful description of the extremes at all angles, the support of the density, $\mathrm{supp}(f_{\mathbf{X}}):=\{\mathbf{x}\in\mathbb{R}^{d}:f_{\mathbf{X}}(\mathbf{x})>0\}$, must be star-shaped with respect to the chosen origin [hansen2020]. That is, given an origin $\mathbf{x}_{0}\in\mathrm{supp}(f_{\mathbf{X}})$ and any point $\mathbf{x}\in\mathrm{supp}(f_{\mathbf{X}})$, the line segment from $\mathbf{x}_{0}$ to $\mathbf{x}$ is contained in $\mathrm{supp}(f_{\mathbf{X}})$. Under this assumption, all rays from the origin reach the ‘edges’ of the distribution without passing through regions which have zero density. In this way, the data-cloud has a well-defined ‘inside’ and ‘outside’, with the ‘outer’ region considered ‘extreme’, and the GP-based representation of the radial component in this region is reasonable. This assumption can be verified by checking plots of the density of observations along various rays from the origin, e.g., histograms of the observed radial variable within small angular ranges. However, as discussed further below, in higher dimensional spaces, a large number of angles are required to obtain a reasonable coverage of the surface of $\mathbb{S}^{d-1}$, and visual inspection of plots of the radial density over each angular range is time consuming. Ultimately a data-driven approach for selecting an optimal choice of origin would be best.
@@
In this example, we use knowledge of the physical processes to define an appropriate origin. Firstly, wave breaking limits the maximum possible wave height for a given wave period, with the limit related to wave steepness given by $s=2\pi H_{s}/(gT_{m}^{2})$. In the three-dimensional subspace containing the variables $(H_{x},H_{y},T_{m})$, this results in a conical-shaped bound on the data, centred along the axis $H_{x}=H_{y}=0$, with the radius of the cone given by $H_{s,max}=s_{max}gT_{m}^{2}/(2\pi)$, where $s_{max}$ is the limiting steepness. The value of $s_{max}$ depends on the water depth and wind speed (among other factors), but the maximum value was found to be around $s_{max}\approx 0.08$ for our dataset. This conical shape to the distribution (see Figures 3 and 7) suggests that an appropriate choice of origin should be somewhere on the axis $H_{x}=H_{y}=0$. Due to the physical dependence of wave height on wind speed, we also locate the origin at $U_{x}=U_{y}=0$. The choice of origin for $L_{T}$ is somewhat arbitrary, but experimentation showed that using the mean value of $L_{T}$ gave satisfactory results.
@@
The angular and radial variables are therefore defined with respect to the normalised variables given by $X_{1}=U_{x}/\mathrm{STD}(U_{x})$, $X_{2}=U_{y}/\mathrm{STD}(U_{y})$, $X_{3}=H_{x}/\mathrm{STD}(H_{x})$, $X_{4}=H_{y}/\mathrm{STD}(H_{y})$, and $X_{5}=(L_{T}-\mathrm{mean}(L_{T}))/\mathrm{STD}(L_{T})$, where $\mathrm{STD}(\cdot)$ denotes the standard deviation function. The pairwise relations between these normalised variables are shown in Figure 3. A radial grid has been overlaid to illustrate that these two-dimensional projections are approximately star-shaped with respect to this origin. Lines of bounding steepness $s=0.08$ are shown in the plots of $(X_{3},X_{5})$ and $(X_{4},X_{5})$ as dashed lines. It can be seen that there is far less scatter in the variables close to these bounds due to the physical limitations. Another feature of the data that is evident is the strong positive correlation between the x- and y-components of the wave height and wind speed (i.e., the pairs $(X_{1},X_{3})$ and $(X_{2},X_{4})$).
@@ 4.3 Exploratory data analysis
Before assessing the model fit, it is useful to consider various visualisations of the data, in order to understand how it is distributed over the five-dimensional space. The plots below the diagonal in Figure 3 show the empirical densities of pairs of angular components $(W_{i},W_{j})$, $i,j\in\{1,...,d\}$. By definition, these variables must fall within the unit circle. Any large gaps in the observed values indicate that it is not possible to fit the model at these angles. The objective of using the MLP model of the angular variation of the radial distribution is to estimate a relatively smooth variation with angle. The model should therefore be able to smooth over small angular ranges with no observations. However, the model is unlikely to be able to accurately estimate the behaviour of the radial component over large angular ranges where there is little or no data.
@@
Figure 4: Empirical joint density of wind direction and wave direction. (Blue = low density, yellow = high density).
To assess the variation in the density of angles around a circle, we could plot the number of observations within discrete angular ranges as a histogram. In higher dimensions, visualisation becomes more difficult, but a similar approach can be taken. We count the number of observations within a small angular range of a pseudo-regularly spaced grid of points on the sphere. It is not possible to define evenly-spaced points on the surface of a (hyper) sphere in three or more dimensions. To address this issue we define a set of points $U$ that are approximately regularly-spaced points on $\mathbb{S}^{d-1}$, using the method described in Appendix A. For each $\mathbf{u}\in U$ we count the number of observed angles with $\arccos(\mathbf{w}_{i}^{\top}\mathbf{u})<\theta_{max}$, where $\theta_{max}$ is some prescribed range. Although there may be some overlap between the ranges defined above, this analysis still gives an indication of how the density of angles varies over the sphere.
Figure 6 shows the empirical cumulative distribution function (CDF) of the number of observations in a cell chosen at random, for a cell radius of $\theta_{max}=15^{\circ}$ and a set of 1002 direction vectors (generated using $m=5$ points along each dimension – see Appendix A). One feature that is evident is that over 50% of angular cells contain no observations. This is due to the particular choice of origin, which was selected to meet the assumption of a star-shaped distribution. Figure 6 also shows the number of observations in each of the $322^{5}=$ 32 orthants in $\mathbb{R}^{5}$ (an orthant is the $d$-dimensional analogue of a quadrant of the plane). There are five orthants which contain no observations, and a further five which contain fewer than 100 observations (out of a total of 271,704 observations). The SPAR model estimates of the extremes in the orthants with little data will therefore be highly uncertain. However, Figure 6 shows scatter plots of the maximum observed values of $H_{s}$ and $U_{10}$ in each of the 1002 angular cells against the corresponding number of observations in the cell, indicating that larger values of these variables tend to coincide with higher angular densities. The higher uncertainties associated with the lower occurrence regions should therefore have less effect on the global extremes.
@@
Figure 7 shows a scatter plot of $T_{m}$ against $H_{x}$ and $H_{y}$. (A plot of the normalised variables $(X_{3},X_{4},X_{5})$ would look similar, but the non-normalised variables are shown here to aid physical interpretation). The conical shaped bound imposed by the limiting wave steepness is evident. Another feature that is apparent is that the data cloud is hollow on the side $H_{x}<0$. This is because of fetch limitations in this direction, meaning that waves propagating towards the west are steeper wind-driven waves, so that $H_{s}$ and $T_{m}$ are strongly correlated in this region. This violates the assumption of the distribution being star-shaped with respect to the choice of origin. However, the ‘edge region’ that is not modelled corresponds to low values of $H_{s}$ at a given $T_{m}$ and wave direction, which is less critical for extreme responses. For the present choice of origin at $(H_{x},H_{y},T_{m})=(0,0,6.2)$, it appears that there are some sharp changes in the distribution with the vertical angle, for rays towards the negative x direction. This might be hard to capture with an ANN. Nevertheless, we can assess how well the model performs, given these challenges.
@@
Figure 7: Scatter plot of observations of x-y components of $H_{s}$ and $T_{m}$. Colours indicate the value of $H_{s}=\sqrt{H_{x}^{2}+H_{y}^{2}}$ (blue=low, yellow=high).
4.4 Assessment of angular density model
The assessments of the models for both the angular and conditional radial distributions are derived from a sample simulated from the fitted model, of 100 times the size of the original sample (i.e., $2.71\times 10^{7}$ samples, or approx. 3000 years of hourly data). The larger sample size reduces the uncertainty in estimates in the lower-probability regions, compared to the observed sample, and allows us to assess how well the model extrapolates from the observations. Figure 8 shows a comparison of the empirical joint densities of paired angular components $(w_{i},w_{j})$, $i,j\in\{1,...,5\}$, with contours of the corresponding joint densities from the KD model. For the empirical density of the observations, the colour resolution has been restricted so that the boundaries of the coloured regions correspond to the contour levels from the KD model. That is, contours from the KD model have been computed at density levels $10^{-3},10^{-2.5},...,10^{-0.5}$, and for the observations, all bins where the estimated density is in the interval, say $[10^{-2.5},10^{-2})$, have been assigned the same colour. Overall, the KD model compares very well with the observed densities, with the contours following the complex patterns in the observations. The smoothing effect of the model on the observations is evident, especially at low density levels.
Figure 8: Comparison of empirical joint densities of pairs of angular components (coloured plots) with contours of the joint densities from a sample from the KD model (red lines). Note that the colour resolution has been restricted so that the boundaries of the coloured regions correspond to the contour levels from the KD model.
Figure 9: Comparison of empirical marginal densities of angular components (histograms) with marginal densities from the KD model (red lines).
Whilst Figure 8 shows that the KD model provides a good representation of dependence between pairs of angles, it does not consider how the model performs locally on small regions of $\mathbb{S}^{4}$. To motivate how this can be assessed in five dimensions, first consider the one-dimensional case. Figure 9 shows the empirical marginal densities of each angular component depicted as histograms, compared with the marginal estimates from the KD model. When we assess the fit of the model in this way, we are making a visual assessment of the whether the number of observations in each bin agrees with the number predicted by the model. This idea can readily be extended to higher dimensions, by partitioning the space into discrete regions and comparing the number of observations in a region, with the expected number in that region from the model. In this case, we use a Voronoi partition relative to a set of pseudo-regularly spaced direction vectors, generated using the method described in Appendix A. A Voronoi partition is a partitioning of a space, where each partition is the set of points closest to a given reference point. Figure 10 shows an example partitioning of the sphere in $\mathbb{R}^{3}$ relative to a set of 102 pseudo-regularly spaced direction vectors.
Suppose that we have a partitioning of $\mathbb{S}^{d-1}$ into non-overlapping cells, together with a model from which we can calculate the probability, $p_{i}$, that an observation chosen at random falls into cell $i$. If the observations are independent, the number of observations, $N_{i}$, falling into cell $i$ follows a binomial distribution with CDF
$\Pr(N_{i}\leq k)=\sum_{j=0}^{k}\binom{n}{j}p_{i}^{j}(1-p_{i})^{n-j},\quad k\in\mathbb{N},$ (15)
where $n$ is the number of independent observations. In our application, the data are serially correlated, and hence violate the above assumption of independence. To mitigate for this, we down-sample the data to take one observation per day, resulting in a sample size of $n=11321$ independent observations.
Equation (15) can be used to add confidence bounds to scatter plots of the observed number in a cell against the expected number predicted by the model. Figure 11 shows a scatter plot of the number of observations in each cell of a Voronoi partition of $\mathbb{S}^{4}$ against the expected number from the KD model. The expected number in each cell is calculated as $\mathrm{E}(N_{i})=np_{i}$, where $n$ is the observed (down-sampled) sample size and $p_{i}$ is the fraction of the simulated sample falling into cell $i$. The Voronoi partition is based on a set of 5890 direction vectors, generated using $m=8$ in the method described in Appendix A. However, only 1001 of the resulting cells contained observations, due to the non-uniformity of the angular distribution, discussed above. Overall, the model does very well at replicating the local characteristics of the density, both in the high and low density regions, with differences attributable to sampling variability.
Figure 10: Example of Voronoi partitioning of the sphere in $\mathbb{R}^{3}$ into non-overlapping cells, corresponding to the set of points closest to each direction vector (black lines).
Figure 11: Comparison of observed number of data points in a local angular cell against the expected number in the cell from the KD model, based on a Voronoi partition of $\mathbb{S}^{4}$. Solid black line shows 1:1 relation, red dashed lines show 95% confidence interval on observed number based on sampling variability.
4.5 Assessment of conditional radial model
We assess the quality of the conditional radial model in several ways, analogous to the diagnostics for the angular model. The ability of the model to recreate the pairwise relations is considered in Figure 12, which shows empirical joint densities of pairs of the observed variables (on the orignal scale) compared with contours of the joint density from the fitted model. Although the SPAR model gives an explicit representation of the joint density, this cannot be projected into lower dimensions without computationally-expensive integration. Instead, the joint densities from the model are estimated empirically from a simulated sample from the model. As such, the density shown is subject to sampling uncertainty in the low density regions, but less than that for the observations, due to the larger simulated sample size. Within the body of the data (i.e., below the threshold level) the observations are resampled rather than modelled. Good agreement between the model contours and observations is therefore to be expected in the bodies of the data clouds. However, the density contours in the outer ‘extreme’ regions correspond to points which are simulated from the non-stationary GP model. Overall, the simulated relationships between variables closely follow those of the observations. The model predicts some larger values of steep waves than observed. However, these occur at lower wave heights, so are less critical for design purposes.